Find the solution of Bernoulli Equation


Given:


2xdy+y(y^2lnx-1)dx=0

1
Expert's answer
2021-12-09T14:12:02-0500

"2xdy+y(y^2lnx-1)dx=0\\
2xdy+(y^3lnx-y)dx=0"

Multiply by "\frac{1}{y^3}" both sides, we get

"\frac{2xdy}{y^3}+(lnx-\frac{1}{y^2})dx=0"

"\begin{aligned}
&\text { Equation in total differentials } M(x, y) \mathrm{d} y+N(x, y) \mathrm{d} x=0 \\
&\text { where } M(x, y)=\frac{2 x}{y^{3}} \text { and } N(x, y)=\ln (x)-\frac{1}{y^{2}} \\
&\text { Check for full differential: } M(x, y)_{x}^{\prime}=N(x, y)_{y}^{\prime}=\frac{2}{y^{3}} \\
&\text { Find } F(x, y): \mathrm{d} F(x, y)=F_{y}^{\prime} \mathrm{d} y+F_{x}^{\prime} \mathrm{d} x \\
&F(x, y)=\int N(x, y) \mathrm{d} x=\int \ln (x)-\frac{1}{y^{2}} \mathrm{~d} x=x \ln (x)-\frac{x}{y^{2}}-x+C_{y} \\
&\qquad\left(x \ln (x)-\frac{x}{y^{2}}-x\right)_{y}^{\prime}=\frac{2 x}{y^{3}} \\
&\qquad \begin{array}{c}
C_{y}=\int M(x, y)-\left(x \ln (x)-\frac{x}{y^{2}}-x\right)_{y}^{\prime} \mathrm{d} y=\int 0 \mathrm{~d} y=0 \\
F(x,y)= x \ln (x)-\frac{x}{y^{2}}-x+C_{y}=x \ln (x)-\frac{x}{y^{2}}-x
\end{array}
\end{aligned}"

"x \ln (x)-\frac{x}{y^{2}}-x=C"